OpenStudy (anonymous):

Solve 12x + 2(5 – 2x) = 2x – 6(7 – x). Please show all of your work.

OpenStudy (anonymous):

12x + 10 - 4x = 2x - 42+ 6x 8x + 10 = 8x -42 8x = 8x -52 0x = -52 x is undefined.

OpenStudy (anonymous):

thanx